The titration of a 10.0ml sample of vinegar requires 30.0ml of 0.20 M NaOH solution. Calculate



i. The molarity


ii. The mass /mass percent concentration of acetic acid.


Given that the density of acetic acid is 1.01g/cm3


1
Expert's answer
2022-04-12T13:10:03-0400

The titration reaction of the acetic acid by the sodium hydroxide:


"CH_3COOH + NaOH = CH_3COONa + H_2O"


i. The molarity of acetic acid titrated can be deduced from the following equation:

"n(CH_3COOH)=n(NaOH)"

"c(CH_3COOH)\times V(CH_3COOH) = c(NaOH) \times V(NaOH)"

"c(CH_3COOH)=\frac{c(NaOH) \times V(NaOH)}{V(CH_3COOH)}=\frac{0.20M \times 30.0 ml}{10.0 ml} = 0.60M"


ii. Mass of the solute is determining by the formula:


"c(CH_3COOH)=\frac{n(CH_3COOH)}{V} = \frac{m \times M}{V}"

"m(CH_3COOH)=\frac{c\times V}{M}=\frac{0.6 mol/L \times 0.01L}{60.05 g/mol}=0.0001g=10^{-5}g"


Mass of the solution:

"m = V\times d = 10.0 ml \times 1.01 g/cm^3 = 10.1 g"


Mass/mass percent concentration of acetic acid:


"W = \frac{m(CH_3COOH)}{m} \times 100 \% = \frac {0.0001 g}{10.1 g}\times 100 = 0.00099 \%"